OpenStudy (anonymous):

a 70 kg persom travels to a planet where he waighs 1,700 newtons what is the value of g on that planet

OpenStudy (s3a):

F = ma F = mg (1700 N) = (70kg)g g = (1700N)/(70kg) g = 24.285714285714285714285714285714 N/kg = 24.285714285714285714285714285714 m/s^2 The following is not part of the answer to your question but is additional information.: (24.285714285714285714285714285714 m/s^2)/(9.8 m/s^2) = 2.4781341107871720116618075801749 which means that the gravity on that other planet is about 2.5 times as strong as that as our planet Earth.

OpenStudy (anonymous):

dem sig figs.

OpenStudy (anonymous):

25